# Can LLMs Recognize Toxicity? Structured Toxicity Investigation Framework and Semantic-Based Metric

### v1

@@ Abstract
In the pursuit of developing Large Language Models (LLMs) that adhere to societal standards, it is imperative to discern the existence of toxicity in the generated text. The majority of existing toxicity metrics rely on encoder models trained on specific toxicity datasets. However, these encoders are susceptible to out-of-distribution (OOD) problems and depend on the definition of toxicity assumed in a dataset. In this paper, we introduce an automatic robust metric grounded on LLMs to distinguish whether model responses are toxic. We start by analyzing the toxicity factors, followed by examining the intrinsic toxic attributes of LLMs to ascertain their suitability as evaluators. Subsequently, we evaluate our metric, LLMs As ToxiciTy Evaluators (LATTE), on evaluation datasets. The empirical results indicate outstanding performance in measuring toxicity, improving upon state-of-the-art metrics by 12 points in F1 score without training procedure. We also show that upstream toxicity has an influence on downstream metrics.
Can LLMs Recognize Toxicity? Structured Toxicity Investigation Framework and Semantic-Based Metric
@@ 1 Introduction
In recent academic literature, Large Language Models (LLMs), such as Llama-variants (Touvron et al., 2023a; Chiang et al., 2023; Touvron et al., 2023b) and GPT-variants (Brown et al., 2020; OpenAI, 2023; Mukherjee et al., 2023), have exhibited excellent performance across a wide variety of downstream tasks. They demonstrate emergent ability to efficiently adapt to various downstream tasks through in-context learning (ICL) framework which yields effects akin to parameter training (von Oswald et al., 2023).
@@
However, LLMs are not free from bias and offensiveness, as bias is deeply ingrained in their training data (Gonen and Goldberg, 2019a). Existing research (Sharma et al., 2021; Roh et al., 2021; Gaci et al., 2022a; Zeng et al., 2022) addresses such issues, focusing on representative identity terms (Sap et al., 2020; Blodgett et al., 2021; Uchida et al., 2022a; Goldfarb-Tarrant et al., 2023). Within this research trajectory, some studies design trigger-prompts reflecting fine-grained toxic categories to assess the toxicity of the LLMs (Gehman et al., 2020; Smith et al., 2022; Deshpande et al., 2023; Wang et al., 2023; Shaikh et al., 2023). The purpose of these methodologies is to assess whether a model exhibits a toxic response when provided with a statement oriented towards toxicity as input.
To determine whether such responses are toxic, typical toxicity metrics such as toxicity classifier scores and PerspectiveAPI [^1] https://perspectiveapi.com/ are deployed. However they are vulnerable to domain shift and perturbations shown in Figure 1 and Appendix A.2. aligned to the results that the use of those approaches show significant susceptibility in the distribution shifts (Pozzobon et al., 2023). It is challenging for fine-tuned classifiers to detect instances that deviate from the definition of toxicity in the dataset (Orgad and Belinkov, 2022; Moradi and Samwald, 2021; Kumar et al., 2022), as more training procedures on datasets introduce catastrophic forgetting.
Another line of work has utilized LLMs to evaluate utterances by themselves and perform self-debiasing to improve policy compliance (Morabito et al., 2023; Qi et al., 2023). However, these methodologies tend to blindly trust LLMs and often overlook the potential adverse impact of upstream bias on the self-debiasing process. Additionally, it remains unclear in what aspects these evaluation methods are better than traditional metrics, and the prompt factors that influence the evaluation performance are under-explored.
In this paper, we propose LATTE (LLMs As ToxiciTy Evaluator) to overcome the aforementioned issues considering two main factors. First, provided that the definition of toxicity can vary dynamically depending on diverse contexts, it is essential to make the metric robust against OOD problems. The metric should function effectively even when toxicity definitions are changed, ideally without having to train on a new dataset. To achieve this, we take advantage of LLM’s zero-shot capabilities. Second, the methodology using foundational neural models should not be indiscriminately applied as value prompts, which are commonly used in reasoning methods, as upstream bias have shown impact downstream tasks significantly (Sun et al., 2022; Feng et al., 2023). Therefore, we aim to identify the safe domains for each LLM by fine-grained toxicity investigation and apply our method only within such domains.
@@
As shown in Figure 2, we first define toxicity factors and set up a benchmark tailored to each factor. Next, we propose a toxicity investigation framework to discover the domains where LLMs are not inherently toxic and maintain neutral corresponding to each factor. Once we find a safe domain, we compare the conventional measurements to LATTE, our proposed evaluation metric. Experimental results reveal that LATTE demonstrates robustness and superior performance by more than 4 points in accuracy and 12 points in F1 score, compared to existing metrics in evaluation datasets. In addition, we provide analysis that our instruction template of evaluation metric is robust to perturbations, and that the neutrality of upstream LLMs does contribute to performances in downstream metrics. Finally, our evaluation prompt is not confined to specific LLMs, but applicable to other LLMs.
2 Definition of Toxicity: Three Factors
In this work, we define toxicity using three distinct factors — demeaning content, partiality, and ethical preference. To articulate the notion of toxicity within our research, we refer to elements of non-maleficence, fairness, and justice from the seven AI trustworthy factors Jobin et al. (2019).
@@ 2 Definition of Toxicity: Three Factors
From the viewpoint of fairness, those that unilaterally support or oppose a specific group or stance can also be perceived as toxic. Smith et al. (2022) deal with fairness based on demographic terms in detail, and Lee et al. (2023a) address argumentative and contentious questions as sensitive questions. In addition, these factors and the demeaning factor do not always coincide with each other (Fleisig et al., 2023), and stereotypes aren’t always negatively assumed (Blodgett et al., 2021). As a result, such elements will be collectively defined as “Partiality”.
Furthermore, the concept of justice is technically interpreted in Hendrycks et al. (2021). From an ethical perspective, individual values are segmented into virtue, deontology, and utilitarianism. LLMs may induce discomfort in users by overlooking individual values, even if the utterances are not explicitly demeaning or biased. We hence categorize such instances as "weak toxic" instances. Those factors will be collectively defined as “Ethical Preference”. From Demeaning to Ethical Preference, it becomes more complicated to reach a consensus on whether the utterances are toxic or not.
Factor Type Prompt Content
Take a deep breath and think step-by-step.
Is the content (A) Demeaning or (B) Respectful ?
{Content} {Question} (A) Not enough info (B) {Identity1} (C) {Identity2}
{Normative Definition} {Controversial Definition}
{Content} Is the content (A) Normative or (B) Controversial ?
Utilitarianism Take a deep breath and think step by step.
@@ 3 Methodology
3.1 Toxicity Investigation
Prior to the implementation of our proposed metric LATTE, it is crucial to ascertain its fairness, as semantic metrics grounded in neural architectures exhibit biases (Sun et al., 2022), and upstream biases have been shown to influence downstream tasks (Feng et al., 2023). We thus investigate the toxicity of LLMs with regards to two aspects: whether the model has an understanding of the concept (Awareness), and whether it also maintains a neutral position with regards to it (Neutrality).
3.1.1 Demeaning
Neutrality Test: Aggression Questionnaire Test (AQ) Inspired by Feng et al. (2023), we adopt the renowned self-report psychology test AQ to LLMs for investigating the extent of neutrality on aggression. AQ defines aggression as the behavior to cause harm to those trying to avoid it (Webster et al., 2014). The test takes four factors into account, namely Physical Aggression, Verbal Aggression, Anger, and Hostility. As Miotto et al. (2022) report that GPT-3 scores similarly to human samples in terms of personality, the average scores of humans from Buss and Perry (1992) are shown in Table 2 for comparison. Details of the scoring procedure are in Appendix D.
3.1.2 Partiality
Neutrality Test: Political Compass Test & Argumentative Test According to the definition of Partiality, responses that favor one side to argumentative or contentious utterances can be problematic. Therefore, we also probe the political and economic orientations of LLMs through the political compass test introduced by Feng et al. (2023). Political compass test assesses LLMs’ political positions in a two-dimensional spectrum. The x-axis represents economic orientation, while the y-axis indicates social orientation in Figure 13.
If the results show that LLMs possess bias towards one side, the LLM is required to undertake additional pretraining or, at least, be able to distinguish such contents. Due to resource constraints on further training LLMs, we instead test whether LLMs can distinguish utterances between argumentative content and demographically biased content as shown in Table 1 Argumentative.
3.1.3 Ethical Preference
@@ 3.2 Toxicity Metric: LATTE
where $\mathbf{x}$ is the input utterance and $\mathbf{y}$ is the label. With the optimal variables ($*{}^{*}$) identified, inference is conducted as follows:
$\begin{split}score_{1}=M^{*}(x_{1}|s^{*},c^{*},i^{*}),\;x_{1}\in D_{test},\\ class_{1}=\left\{\begin{array}[]{cl}1&:\ score_{1}\geq t,\\ 0&:\ score_{1}<t,\end{array}\right.\end{split}$ (1)
where $D_{test}$ is the test dataset and $t$ is the threshold. Once the inherent toxicity elements of LLMs ($M$) in Equation 1 are thoroughly investigated in Section section 3.1, we can then qualify or disqualify them to act as toxicity evaluators ($M^{*}$).
@@
Content $c$ As the Chain-of-Thought (CoT) prompting methodology significantly influences the performance of diverse NLP tasks (Kojima et al., 2023; Shaikh et al., 2023; Yang et al., 2023), we consider such a reasoning method in our evaluation prompt. In addition, we append words that have the same meaning, but in a different language as controlling non-target languages in multilingual LLMs has a substantial effect on the overall gender bias performance (Lee et al., 2023b). Furthermore, in the semantic dimension, the effects of adding an antonym of the word and the definition of toxicity are also tested to prevent potential variations in performance.
Interval $i$ Lastly, there are choices related to the scoring scale — 0 to 1, 1 to 10, and 1 to 100. These factors are also reflected in the evaluation prompt. All examples are in Appendix B.
Due to the characteristics of toxicity that vary depending on the dataset, we utilize the same prompt during the evaluation stage, except for the definition of toxicity. Next, we apply our evaluation prompt into other LLMs to demonstrate its generalizability.
@@ 4 Experiment
In Section section 4.1, we introduce datasets for investigating the model’s inherent toxicity. In Section section 4.2, we set up datasets for experimenting the feasibility of LLMs being used as evaluators. Former datasets are primarily designed for detecting toxic utterances, whereas the purpose of toxicity detection is auxiliary for latter datasets. All the examples of investigation dataset are in Appendix C. For metrics, we utilize the task performance accuracy and F1 score.
4.1 Toxicity Investigation
@@ 4.1.1 Demeaning Datasets
FairPrism (Fleisig et al., 2023) is an English dataset covering a diverse set of harms, containing context-dependent harms. In this work, we utilize the demeaning category. HateSpeech (Yoder et al., 2022) is a dataset focusing on English texts, and categorizes data into demographic, identity, and hegemony subsets.
@@ 4.1.2 Partiality Datasets
BBQ (Parrish et al., 2022) is constructed for the Question Answering (QA) task, and is comprised of 11 stereotype categories. Followed by sampling 100 examples from each group, a total of 1,100 samples are converted into a multiple choice-QA format to test LLMs’ demographic bias. SQUARE Demographic (Lee et al., 2023a) comprises of sensitive questions and responses based on Korean culture. In our work, we obtain test sets from inclusive-opinion and ethically-aware categories, for measuring the awareness of demographic content and argumentative contents.
4.1.3 Ethical Preference Dataset
ETHICS (Hendrycks et al., 2021) utilizes natural language scenarios to generate numerous situations, encompassing interpersonal dynamics and daily events. In our experimental design, we use the deontology, virtue, utilitarianism test scenarios.
@@ 4.1.3 Ethical Preference Dataset
4.2 Toxicity Metric : LATTE
4.2.1 Evaluation Datasets
ParaDetox (Logacheva et al., 2022) is a short text paraphrased dataset of toxic and neutral utterances, filtered from the Jigsaw, Reddit, and Twitter datasets. They define toxicity as the use of profanity. In our definition, it covers the demeaning factors. In our work, we utilize toxic and the neutral utterances. Prosocial Dialogue (Kim et al., 2022) is a conversational dataset incorporating prosocial norms. We define Need-Caution utterances as toxic utterances and casual utterances as non-toxic utterances. In our definition, it covers both demeaning factor and demographic-oriented partiality factor. SQUARE Contentious Other than SQUARE Demographic, we use contentious and unacceptable utterances for investigating how neutrality is important in the downstream stage. All the details of evaluation datasets are in Appendix G.
4.3 Evaluation Baselines
[…]
4.4 LLMs
Our target models are GPT3-text-davinci-003, GPT-3.5 turbo, GPT-4, and Llama2 7B, 13B, 70B. We use 4 A6000 gpus for Llama2 and OpenAI’s API for GPT333versions prior to January 20, 2024. In the toxicity metric evaluation stage, we use the LLMs qualified by our proposed standards with our proposed evaluation prompt. We utilize deterministic decoding strategies to eliminate randomness in the measurement and to guarantee consistent agreement score for a fixed input text, except the case where LLMs need to generate sentences. We adopt non-deterministic decoding with default parameters for the task of generating sentences for evaluations such as the political compass test.
5 Results & Analysis
5.1 Toxicity Investigation
5.1.1 Demeaning
@@
In Figure 3, the red baseline represents the performance of SBERT trained on the original training data. For LLMs, models with a smaller number of parameters tend to underperform, and their performances are notably sensitive to changes in the prompt. We empirically discover that interference also occurs in smaller Llama2 models. Such observations align with the claims that multilingual LLMs should guarantee a size of parameters proportional to the size of training data (Shaham et al., 2023).
Though the models are able to discern demeaning contents, both GPT-3 and GPT-3.5 show high aggression scores in Table 2. Llama2 70B and GPT-4, on the other hand, are closer to the human, and Llama2 70B yields less aggressive scores compared to GPT-4. We hypothesize that scores of Llama2 70B are less assertive due to the incorporation of a safety module during its training stage.

### v3

# Can LLMs Recognize Toxicity? Definition-Based Toxicity Metric

@@ Abstract
In the pursuit of developing Large Language Models (LLMs) that adhere to societal standards, it is imperative to detect the toxicity in the generated text. The majority of existing toxicity metrics rely on encoder models trained on specific toxicity datasets, which are susceptible to out-of-distribution (OOD) problems and depend on the dataset’s definition of toxicity. In this paper, we introduce a robust metric grounded on LLMs to flexibly measure toxicity according to the given definition. We first analyze the toxicity factors, followed by an examination of the intrinsic toxic attributes of LLMs to ascertain their suitability as evaluators. Finally, we evaluate the performance of our metric with detailed analysis. Our empirical results demonstrate outstanding performance in measuring toxicity within verified factors, improving on conventional metrics by 12 points in the F1 score. Our findings also indicate that upstream toxicity significantly influences downstream metrics, suggesting that LLMs are unsuitable for toxicity evaluations within unverified factors.
Can LLMs Recognize Toxicity? Definition-Based Toxicity Metric
@@ 1 Introduction
Large Language Models (LLMs) are prone to exhibiting biases and generating offensive content, as the bias deeply ingrained in their training data (Gonen and Goldberg, 2019a). Existing research (Sharma et al., 2021; Roh et al., 2021; Gaci et al., 2022a; Zeng et al., 2022) addresses such issues, focusing on representative identity terms or trigger-prompts reflecting fine-grained toxic categories (Sap et al., 2020; Gehman et al., 2020; Shaikh et al., 2023). These studies utilize typical toxicity metrics such as toxicity classifier scores and PerspectiveAPI [^1] https://perspectiveapi.com/ to determine whether a LLM’s reaction is toxic. However, typical metrics are vulnerable to domain shift and perturbations shown in Figure 1 and Appendix A.2. Such phenomenon aligns with the results that the use of those approaches shows significant susceptibility in the distribution shifts (Pozzobon et al., 2023). As those metrics are trained on a particular dataset, they struggle to identify instances that deviate from the predefined notion of toxicity within the dataset (OOD of Toxicity) (Orgad and Belinkov, 2022; Moradi and Samwald, 2021; Kumar et al., 2022).
@@
Recently some researchers adopt LLMs to evaluate utterances by themselves and make LLMs self-debiasing to improve policy compliance (Morabito et al., 2023; Qi et al., 2023). However, these methodologies tend to blindly trust LLMs and often overlook the adverse impact of upstream bias on the self-debiasing process. Additionally, it remains unclear in what aspects these evaluation methods are better than traditional metrics, and what prompt factors influence the evaluation performance.
In this paper, we propose LATTE (LLMs As ToxiciTy Evaluator) to address the variability in definitions of toxicity depending on context and to mitigate the negative impact of upstream bias. First, provided that the definition of toxicity can vary dynamically depending on diverse contexts, it is essential to ensure that the metric is flexible enough to adapt to diverse contexts. That is, the metric should function effectively even when toxicity definitions change, ideally without having to train on a new dataset. To achieve this, we take advantage of LLM’s zero-shot capabilities with our proposed evaluation prompt. Second, the methodology using foundational neural models should not be indiscriminately applied as value prompts, which are commonly used in reasoning methods, as upstream bias has a significant impact on downstream tasks (Sun et al., 2022; Feng et al., 2023). Therefore, we identify the safe domains for each LLM by a structured toxicity investigation, and apply our method only within such domains.
@@
As shown in Figure 2, we first define toxicity factors and set up a benchmark tailored to each factor. Next, we propose a toxicity investigation process incorporating the concept of neutrality to identify the safe domains where LLMs are not inherently toxic and maintain a neutral stance with respect to each factor. When LLMs are used as toxicity evaluators in unsafe domains, we demonstrate that their evaluation performance cannot be trusted. Therefore, it is essential to identify safe domains. Once finding a safe domain, we compare the conventional measurements to LATTE, our proposed evaluation metric. Experimental results reveal that LATTE demonstrates superior performance by more than 4 points in accuracy and 12 points in F1 score, compared to existing metrics in evaluation datasets. In addition, our metric are robust to changes in definition of toxicity and format perturbations. Lastly, we show that the neutrality of upstream LLMs does contribute to performances in downstream metrics, and our LATTE approach can be adaptable to diverse LLMs.
2 Definition of Toxicity
In this work, we define toxicity using demeaning content and partiality factors. To articulate the notion of toxicity within our research, we refer to elements of non-maleficence, fairness.
@@ 2 Definition of Toxicity
From the viewpoint of fairness, language that unilaterally oppose a specific group or stance can also be perceived as toxic. Smith et al. (2022) deal with fairness based on demographic terms in detail, and Lee et al. (2023a) address argumentative and contentious questions as sensitive questions. In addition, such a factor and the demeaning factor do not always coincide with each other (Fleisig et al., 2023), and stereotypes aren’t always negatively assumed (Blodgett et al., 2021). As a result, such elements are collectively defined as Partiality.
@@ 3 Methodology
3.1 LLM Sanity Check : Toxicity
Prior to the implementation of our proposed metric LATTE, it is crucial to ascertain its fairness, as semantic metrics grounded in neural architectures exhibit biases (Sun et al., 2022), and upstream biases have an influence on downstream tasks (Feng et al., 2023). We thus investigate the toxicity of LLMs with regards to two aspects: whether the model has an understanding of the concept (Awareness), and whether it also maintains a neutral position with regards to toxicity factors (Neutrality).
3.1.1 Awareness
To assess the capability of LLMs to detect toxic contents, we propose the investigation framework to discern the awareness of each toxicity factor by designing a task within a zero-shot setting equipped with investigation prompt. For Demeaning, we test the capability of detecting profanities or insults. For Partiality, we measure awareness of partiality based on identity terms.
3.1.2 Neutrality
As Miotto et al. (2022) report that GPT-3 responses are similar to human responses in terms of personality, we adopt the renowned self-report psychology test AQ (Webster et al., 2014) to LLMs for investigating the extent of neutrality on aggression for Demeaning.
Next, responses that favor one side to argumentative or contentious utterances can be problematic. Therefore, we also probe the political and economic orientations of LLMs through the political compass test introduced by Feng et al. (2023) for Partiality. If the results show that LLMs possess bias towards one side, the LLM is required to undertake additional pretraining or, at least, be able to distinguish such contents. Due to resource constraints on further training LLMs, we instead test whether LLMs can distinguish utterances between argumentative content and demographically biased content. All the details are in Appendix B.
@@ 3.2 Toxicity Metric: LATTE
where $\mathbf{x}$ as the input utterance and $\mathbf{y}$ as the label. With the optimal variables () identified, inference is conducted as follows :
$\begin{split}&score_{1}=M^{*}(x_{1}|s^{*},c^{*},i^{*},\text{def}),\;x_{1}\in D_{test}\\ &class_{1}=\left\{\begin{array}[]{cl}1&:\ score_{1}\geq t\\ 0&:\ score_{1}<t\end{array}\right.\end{split}$ (1)
$D_{test}$ is the test dataset and $t$ as the threshold. Once the inherent toxicity elements of LLMs ($M$) in Equation 1 are thoroughly investigated in Section section 3.1, we can then qualify or disqualify them to act as toxicity evaluators ($M^{*}$).
Definition def We provide the definition of toxicity based on each dataset.
@@
Content $c$ As the Chain-of-Thought (CoT) prompting methodology significantly influences the performance of diverse NLP tasks (Kojima et al., 2023; Shaikh et al., 2023; Yang et al., 2023), we consider such a reasoning method in our evaluation prompt. In addition, we append words that have the same meaning, but in a different language as controlling non-target languages in multilingual LLMs has a substantial effect on the overall gender bias performance (Lee et al., 2023b). Furthermore, in the semantic dimension, the effects of adding an antonym of the word.
Interval $i$ Lastly, there are choices related to the scoring scale — 0 to 1, 1 to 10, and 1 to 100. These factors are also reflected in the evaluation prompt. All examples are in Appendix D.
Due to the characteristics of toxicity that vary depending on the dataset, we utilize the same prompt during evaluation, except for the definition of toxicity. Next, we apply our evaluation prompt into other LLMs to demonstrate its generalizability.
@@ 4 Experiment
In Section section 4.1, we introduce datasets for investigating the model’s inherent toxicity. In Section section 4.2, we set up datasets for experimenting the feasibility of LLMs being used as evaluators. Former datasets are primarily designed for detecting toxic utterances, whereas the purpose of detecting toxicity is auxiliary for latter datasets. All the examples of investigation dataset are in Appendix C. For metrics, we utilize the task performance accuracy and F1 score.
4.1 LLM Sanity Check : Toxicity
@@ 4.1.1 Demeaning Datasets
FairPrism (Fleisig et al., 2023) is a representative English dataset covering a diverse set of harms, containing context-dependent harms, enhancements to existing demeaning datasets such as RealToxicityPrompts (Gehman et al., 2020), BOLD (Dhamala et al., 2021), ToxiGen (Hartvigsen et al., 2022). In this work, we utilize the demeaning category. HateSpeech (Yoder et al., 2022) is a dataset focusing on English texts, analyzing an incitement of emotion and violence in hate speech instances, reflecting numerous toxicity datasets such as Civil Comments (Borkan et al., 2019), Social Bias Inference Corpus (Sap et al., 2020), and Contextual Abuse Dataset (Vidgen et al., 2021).
@@ 4.1.2 Partiality Datasets
BBQ (Parrish et al., 2022) is constructed for the Question Answering (QA) task, and is comprised of 11 stereotype categories. Followed by sampling 100 examples from each group, a total of 1,100 samples are converted into a multiple choice-QA format to test LLMs’ demographic bias. SQUARE (Lee et al., 2023a) comprises of sensitive questions and responses based on Korean culture. In our work, we obtain test sets from inclusive-opinion and ethically-aware categories, for measuring the awareness of demographic content and argumentative contents. In our work, we designate those test sets as SQUARE Demographic.
4.2 Toxicity Metric : LATTE
4.2.1 Evaluation Datasets
ParaDetox (Logacheva et al., 2022) is a short text paraphrased dataset of toxic and neutral utterances, filtered from the Jigsaw, Reddit, and Twitter datasets. They define toxicity as the use of profanity. In our definition, it covers the demeaning factors. We utilize toxic and the neutral utterances. Prosocial Dialogue (Kim et al., 2022) is a conversational dataset incorporating prosocial norms. We define Need-Caution utterances as toxic utterances and Casual utterances as non-toxic utterances. In our definition, it covers both demeaning factor and demographic-oriented partiality factor. SQUARE Contentious Other than SQUARE Demographic, we use contentious-unacceptable utterance pairs for investigating how neutrality is important in the downstream stage. All the details of evaluation datasets are in Appendix E.
4.3 Evaluation Baselines
[…]
4.4 LLMs
Our target models are GPT-3-text-davinci-003, GPT-3.5 turbo, GPT-4, GPT-4o and Llama2 7B, 13B, 70B, Llama3 70B. We use 4 A6000 gpus for Llama2 and OpenAI’s API for GPT333versions prior to January 20, 2024. In the toxicity metric evaluation stage, we use the qualified LLMs with our proposed evaluation prompt. We utilize deterministic decoding strategies to eliminate randomness in the measurement and to guarantee consistent agreement score for a fixed input text, except the case where LLMs need to generate sentences. We adopt non-deterministic decoding with default parameters for the task of generating sentences such as the political compass test.
5 Results & Analysis
5.1 LLM Sanity Check : Toxicity
5.1.1 Demeaning
@@
In Figure 3, the red baseline represents the performance of SBERT trained on the original training data. For LLMs, models with a smaller number of parameters tend to underperform, and their performances are notably sensitive to changes in the prompt. We empirically discover that interference also occurs in smaller Llama2 models. Such observations are aligned with the claims that multilingual LLMs should guarantee a size of parameters proportional to the size of training data (Shaham et al., 2023).
Though the models are able to discern demeaning contents, both GPT-3 and GPT-3.5 show high aggression scores in Table 1. Llama2 70B and GPT-4, on the other hand, are closer to the human averages, and Llama2 70B yields less aggressive scores compared to GPT-4. We hypothesize that scores of Llama2 70B are less assertive due to the incorporation of a safety module during its training stage. All the detailed results are in Appendix B.1.1